OpenStudy (anonymous):

how do you solve 3squareroot of x^3y?

OpenStudy (anonymous):

\[=\sqrt[3]{x ^{3y}}\] is this your expression?

OpenStudy (anonymous):

\[=(x ^{3y})^{1/3}=x ^{3y*1/3}=x ^{y}\]

OpenStudy (anonymous):

sorry its the cubed root of x^3y

OpenStudy (anonymous):

what I guessed... see the above :)

OpenStudy (anonymous):

so the answer is x ^y

OpenStudy (anonymous):

or x cubed root of y

OpenStudy (anonymous):

not : x cubed root of y"... that's is incorrect statement. but "answer is x ^y" or x in power of y is correct one.

OpenStudy (anonymous):

cubed root means x in power of (1/3) like square root is power of (1/2). do you understand the difference....?